Consider this probability mass function: px) - 2k forx- 2, 4, 6, 8, 10, 12 Determine k. Enter your answer to 4 decimal places using conventional rounding methods
Answer #1

The sum of all the probabilities should be 1. Hence,

P(2) + P(4) + P(6) + P(8) + P(10) + P(12) = 1

2k + 2k + 2k + 2k + 2k + 2k = 1

12k = 1

k = 1/12

k = 0.0833